OpenStudy (anonymous):

18. For which system of equations would elimination be the simplest technique to choose? (Points : 2) y = 2x and x + 3y = 5 4x + 3y = 6 and 3y = 5 x + y = 7 and x - y = 12 3x - 2y = 8 and 4x - 5y = 20 19. Substitution would be the simplest method of solution for which system? (Points : 2) 5/3 x = 6y and 2x -3y = 8 y = 2x and x + y = 28 3x - 2y =18 and 4x + 5y =60 20. The sum of two numbers is 110. Their difference is 28. What is the smallest of the two numbers? (Points : 2) 56 41

OpenStudy (anonymous):

Elimination would be simplest for: 3x - 2y = 8 and 4x - 5y = 20 y = 2x and x + 3y = 5 ---> You already have what y equals, so substitution is easier. 4x + 3y = 6 and 3y = 5 ---> You can easily solve for y ( y = 5/3) Therefore, substitution is easier. x + y = 7 and x - y = 12 ---> There are no coefficients so solving for either x or y would be simple, then you could easily substitute.

OpenStudy (anonymous):

O ok thanks :)